OpenStudy (derrickvhs):

Find the x-intercepts of the parabola with vertex (-4,2) and y-intercept (0,-30). Write your answer in this form: (x1,y1),(y1,y2). If necessary, round to the nearest hundredth.

OpenStudy (arnavguddu):

assuming that this parabola is inclined....etc.... so use the general equation Ax^2 + 2B xy + Cy^2 + Dx + Ey + F = 0 as these points lie on the parabola..... put them and solve.... shortcut.... you need x intercept where y = 0 put y=0 in this eqn - Ax^2 + 0 + 0 + Dx +0 + F = 0 Ax^2 + Dx + F = 0 Now eqn is reduced to x only, where again we have one point (0,-30)....here put x=0 A*0 + D*0 + F = 0 => F=0 eqn is now Ax^2 + Dx = 0 now with point (-4,2), x = -4 A (-4)^2 + D(-4) = 0 A*16-4D = 0 16A = 4 D 4A = D or D = 4A now put this in eqn Ax^2 + Dx = 0 => Ax^2 + 4Ax = 0 =>x^2 + 4x =0 =>x(x+4) = 0 so x =0, x=-4 are intercepts on x axis for this parabola